A New Jersey motorcycle manufacturer (Hot Sizzle) has the option to make or buy tires. The annual requirement is 22,000 tires. A supplier is able to supply the tires for $80 each. Hot Sizzle estimates that it costs $500 to prepare the contract with the supplier. To make the parts in-house, Hot Sizzle must invest $20,000 in capital equipment and estimates that the tires cost $50 each.

Assuming that cost is the only criterion, use break-even analysis to determine whether Hot Sizzle should make or buy the item. What is the break-even quantity?

500

1000

650

750

900
Answer #1

650

Let the break even point for the processes be X, then

500 + 80*X = 20000 + 50*X

30*X = 19500

X = 650 Units